1. The number of radians the minute hand will rotate in 35 minutes is...?
2. the reference angle for 222.2 degrees is....?
3. Use identities to find tan A if sec A = 7/3 and A is in quadrant IV?
4. The signs of sec 444 degrees and tan 444 degrees are....?

Answer by stanbon(75887)

1. The number of radians the minute hand will rotate in 35 minutes is...?
Ans: (35/60)2pi = 3.665 radians
------------------------------------------------
2. the reference angle for 222.2 degrees is....?
Ans: 222.2-180 = 42.2 degrees
-----------------------------------
3. Use identities to find tan A if sec A = 7/3 and A is in quadrant IV?
sec = r/x, so r = 7 and x = 3
----
Note:: y is negative in QIV
----
Then y = -sqrt[7^2-3^2] = -sqrt(40) = -2sqrt(10)
Ans: tan(A) = y/x = -2sqrt(10)/3
===========================================
4. The signs of sec 444 degrees and tan 444 degrees are....?
444-360 = 84 degrees
sec(444) = sec(84) is positive in QI
tan(444) = tan(84) is positive in QI
================
Cheers,
Stan H.
----------------